OpenStudy (anonymous):

Sonia purchased a pool for $2,150 using a six-month deferred payment plan. The interest rate after the introductory period is 27.99%. No down payment is required, but there is a minimum monthly payment of $35. What is the balance at the beginning of the seventh month if only the minimum payment is made each month during the introductory period? help plzz

ganeshie8 (ganeshie8):

lol they don't even change the numbers in problems is it. Exact same problem here : https://answers.yahoo.com/question/index?qid=20120225133428AAghrEl

OpenStudy (anonymous):

This webpage is not available

OpenStudy (igreen):

Oh, well the answer is here: 35 * 6= 210 in payments made A= P( 1 +r/n)^(n*t) A = 2150( 1 + 0.2799/12)(12*.5) = 2468.99 amount deferred 2468.99 - 210 paid = $2258.99 Most deferred payment plans charge the interest from date of purchase, it is NOT applied to the loan as long as it is paid off before the end of the deferred period.

OpenStudy (anonymous):

thnx

OpenStudy (uchezzy):

i did the saame thing

OpenStudy (anonymous):

thnx guys

OpenStudy (igreen):

No problem.

OpenStudy (uchezzy):

anytime u need help

OpenStudy (anonymous):

Karolyn bought a water softener from a rental center for $1,850. She makes 16 monthly payments of $162.75 with her credit card. The rental center charges a $2.25 fee for every late payment made with a credit card. What is the total cost of the water softener if each payment is late? Could you guys help me with this i did it $2,604

OpenStudy (uchezzy):

the equation i think 16 x $162.75 Then add 16 x $2.25 (late fee)

OpenStudy (anonymous):

wait the answer is 2.604

OpenStudy (uchezzy):

i think idk

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@iGreen help me

OpenStudy (uchezzy):

was that wrong

OpenStudy (anonymous):

i think itswrong cause we don't use $1,850

OpenStudy (uchezzy):

hmmmmmmmmmmmmmm

Miracrown (miracrown):

16 x $162.75 Then add 16 x $2.25 (late fee) :)

OpenStudy (anonymous):

its 2604 thats the answer but what should i do with 1850

OpenStudy (uchezzy):

told u

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

it's wrong @Miracrown @uChezzy

OpenStudy (anonymous):

the answer is wrong i got 0

Miracrown (miracrown):

Did you added like I told you to?

OpenStudy (anonymous):

yes

Miracrown (miracrown):

What did you write as an answer?

OpenStudy (anonymous):

they change thequestion

OpenStudy (anonymous):

2604

Miracrown (miracrown):

No, see, you did it wrong. You didn't even bother following my instructions. I told you to, calculate this first: 16 x $162.75 which = 2604 and then, I told you to: add 16 x $2.25 which = 36 So, 2604+36 = 2640, my friend. :)

OpenStudy (anonymous):

oh :/

OpenStudy (uchezzy):

bossy

Miracrown (miracrown):

Yesh, don't rush with your work!

Miracrown (miracrown):

Don't worry, we all make mistakes - nothing to be discouraged about.

OpenStudy (anonymous):

thank you :/ help me with this please Karen bought a dishwasher from a rental center for $650. She makes 16 monthly payments of $63.75 with her credit card. The rental center charges $1.50 for every payment made with a credit card. What is the total cost of the dishwasher?

OpenStudy (uchezzy):

especially me

OpenStudy (uchezzy):

$1,044 i google it

OpenStudy (uchezzy):

its from this site so

OpenStudy (anonymous):

are the google answers are right most of the time i google the answers i always got them wrong

Miracrown (miracrown):

They aren't reliable, just give it a shot at it yourself and then if you have any doubts with it, don't hesitate to ask any one of us. We're all here to help you! :)

OpenStudy (uchezzy):

yea what miracrown said

OpenStudy (anonymous):

Sarah purchased a swing set for $1,450 using a six-month deferred payment plan. The interest rate after the introductory period is 19.99%. No down payment is required, but there is a minimum monthly payment of $20. What is the balance at the beginning of the seventh month if only the minimum payment is made each month during the introductory period? what about this one &*_*